OpenStudy (kl0723):

Determine if the series is convergent or divergent, if it's convergent, find the sum:

OpenStudy (kl0723):

\[\sum_{n=1}^{\infty}\frac{ (-3)^{n-1} }{ 4^n }\] can u guys help me out ?

OpenStudy (anonymous):

Recall that for \(|r|<1\), \[\sum_{n=0}^\infty ar^n=\sum_{n=1}^\infty ar^{n-1}=\frac{a}{1-r}\] In this case, \(r=-\dfrac{3}{4}\), so the series indeed converges. \[\frac{(-3)^{n-1}}{4^n}=\frac{(-3)^{n-1}}{4\times4^{n-1}}=\frac{1}{4}\left(-\frac{3}{4}\right)^{n-1}\] so \(a=\dfrac{1}{4}\).

OpenStudy (kl0723):

how do you get to express the 4^n to (4x4^n)?

OpenStudy (kl0723):

I meant 4x4^(n-1)

OpenStudy (kl0723):

 I'm guessing like this